2. Here some raw data on recombination between Neurospora genes black spores (bla), curved beaks (cur), purple hyphae (pur), spermidine (spe), ropy growth (ro). Based on these data, map the chromosome based on linkage between these five genes. Hint: Determinationsfor short distance are more accurate than determination for longer distance. Extra point: Why is that? # of progen # of recombinants between the genes Genes in crOSS bla, cur cur, ro bla, spe bla, ro cur, pur pur, spe cur, spe pur, ro spe, ro bla, pur 62,679 1,720 685 20,153 51,136 11,985 10,042 13,601 2,054 48,931 14,237 141 326 3,578 10,205 5,474 3,037 1,609 738 3,026
Answer #1

Genetic map:

Recombination frequency (Number of recombinants between genes+ total number of progeny)*100

Using above equation recombination frequency for all different combinations of genes is calculated.

Bla (A), Cur (B)= 22.71%

Cur (B), Ro (C)=8.19%

Bla (A), Spe (D)=47.59%

Bla (A), Ro (C)=17.75%

Cur (B), Pur (E)=19.95%

Pur (E), Spe (D)=45.67%

Cur (B), Spe (D)=30.24%

Pur (E), Ro(C)=11.83%

Spe (D), Ro(C)=35.92%

Bla (A), Pur (E)=6.18%

Recombination frequency indicates rough estimate between the genes. Larger the recombinant frequency larger is the distance between the genes.

Based on this,

A-B=22.71

A-C=17.75

A-D=47.59

A-E=6.18

B-C=8.19

B-D=30.24

B-E=19.95

C-D=35.92

C-E=11.83

D-E=45.67

6.18 11.83 8.19 A B 30.24